How many ounces of a 12% alcohol solution and a 30% alcohol solution must be combined to obtain 90 ounces of a 25% solution?
Answer by addingup(3677)

0.12x+0.30(90-x) = 0.25(90)
0.12x+27-0.30x = 22.5
-0.18x = -4.5
x = -4.5/-0.18
x = 25
You need 25oz of the 12% and 90-25 = 65oz of the 30% to get 90oz at 25%